OpenStudy (anonymous):

What is the area of the triangle? A. 80 mm² B. 160 mm² C. 3200 mm² D. 6400 mm² drawing a pic below

OpenStudy (anonymous):



OpenStudy (anonymous):

@ooxBRITTANYxoo @xMissAlyCatx

OpenStudy (anonymous):

b

OpenStudy (anonymous):

are u 100 prcent sure @xav

OpenStudy (anonymous):

yH

OpenStudy (xmissalycatx):

To find the area of a triangle you use the formula A= 1/2 b(h) Which means a= 1/2 (80) (80) Just take 1/2 of 80 which is 40, then multiply it by 80.

OpenStudy (anonymous):

its 3200

OpenStudy (thread252):

80*80=6400 6400/2=3200 area of triangle = 3200

OpenStudy (anonymous):

thats what i did

OpenStudy (xmissalycatx):

Wow, why didn't I get the medal? XD

OpenStudy (anonymous):

there you go medal